OpenStudy (anonymous):

Sharita does special typing jobs for scientists. She charges $13 for the first 3 pages and $3 for each additional page of scientific typing. Sonia, a biologist, paid $103 to have a research paper typed. How many pages long was Sonia's paper?

OpenStudy (anonymous):

133(n-3)=103 13+3n-9=103 3n+4=103 3n=99 n=33

OpenStudy (anonymous):

23